What is the threshold frequency (in s−1) for cobalt metal if a photon with frequency 1.32 ✕ 1015 s−1 ejects a photon with 7.18 ✕ 10−20 J kinetic energy?
Answer #1

Energy of photon = plank constant * frequency

= 6.626*10^-34 * 1.32*10^15

= 8.74632*10^-19 joules

Total energy = threshold energy + kinetic energy

= 8.75*10^-19 - ( 7.18*10^-20) = KE

= 8.032*10^-19 Joules

Use again energy formula

= 8.032*10^-19 = 6.626*10^-34 * frequency

Frequency = 1.21*10^15 s-1

If you have any query please comment

If you satisfied with the solution please like it thankxx